# WorldPrompter: Traversable Text-to-Scene Generation

### v1

@@ 4.3 Ablation Study
4.3.1 Mixed Training Data
We conduct an experiment by excluding the additional image panorama data. As shown in Fig. 6, incorporating image data significantly enhances the generalization capability of the video generator across a wide variety of scenes, and improves visual quality. Notably, due to the masked diffusion loss training, ground-truth pixels in the bottom regions of the panoramic videos are unavailable. The image data compensate for this limitation, enabling the model to intelligently auto-inpaint the bottom regions, resulting in a natural and realistic appearance.
4.3.2 Masked Diffusion Loss
@@
A cozy cabin library overlooking a vast snowy canyon.
360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
Ours
A coastal castle perched on a cliff under a dramatic sunset.
360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Figure 5: Comparison of 360° video generation results. We present a comparison of 360° panoramic video generation results against 360DVD [42]. Our generated videos exhibit significantly better visual quality and prompt alignment. We manually mark an anchor region with a green box in the 360° image, highlighting the substantially longer traversal range of our generated videos.
A cozy cottage surrounded by wildflowers that glow in the moonlight, nestled deep in an enchanted forest.
w/o Image Data Refer to caption Refer to caption
w/ Image Data
A serene desert where the sand shifts in swirling patterns, reflecting the colors of the sunset.
w/o Image Data Refer to caption Refer to caption
w/ Image Data
Figure 6: We verify that incorporating additional panorama image data helps the model generalize to diverse environments. Without the image data, the generated 360° videos not only exhibit degraded visual appearance but also display messy and incorrect content at the bottom of the video. This issue arises because these pixels lack supervision due to the masked loss.
A garden where every flower blooms in perfect circles, their petals reflecting the colors of the stars above.
Baseline Refer to caption Refer to caption
w/ Masked Loss
A room where the ceiling is made of glass, revealing the swirling, ever-changing night sky beyond.
Baseline Refer to caption Refer to caption
w/ Masked Loss
Figure 7: We conduct an ablation study to evaluate the impact of the masked diffusion loss. This loss function effectively guides the model to generate only the pixels of the scene, eliminating the moving photographer present in the original video training data. As a result, it enables the generation of a complete 360° scene capture for reconstruction.
5 Conclusion
In conclusion, we presented WorldPrompter, the first generative pipeline capable of producing a navigable 3D scenes from text prompts. By leveraging 360° panoramic video generation as an intermediate representation and a fast 3D reconstruction model, WorldPrompter enables the synthesis of immersive environments with comprehensive view coverage and high visual quality. This approach bridges the gap between text-to-3D generation and full scene synthesis.
Although WorldPrompter achieves state-of-the-art performance, generating longer panoramic videos, spanning minutes rather than seconds, remains an open challenge. Consistent content creation over extended sequences is crucial for scaling up to larger, more detailed 3D scenes. Addressing this challenge would further expand applications in virtual worlds and beyond, paving the way for future advancements in scene-level generative models.

### v4

@@ 4.3. Ablation Study
4.3.1. Mixed Training Data
We conduct an experiment by excluding the additional image panorama data. As shown in Fig. 5, incorporating image data significantly enhances the generalization capability of the video generator across a wide variety of scenes, and improves visual quality. Notably, due to the masked diffusion loss training, ground-truth pixels in the bottom regions of the panoramic videos are unavailable. The image data compensates for this limitation, enabling the model to intelligently auto-inpaint the bottom regions, resulting in a natural and realistic appearance.
4.3.2. Masked Diffusion Loss
@@
4.4. Discussions and Limitations
The quality of our text-to-scene generation pipeline is constrained by both the length and resolution of the panoramic video generative model. On one hand, generating minute-long panoramic videos while maintaining visual consistency across extended sequences is essential for scaling to larger, more detailed 3D scenes. On the other hand, panoramic videos allocate more pixels to capture global 3D content and ensure spatial consistency, but this comes at the cost of fine-grained details compared to perspective videos at the same resolution. We address this limitation with an upsampler, though a more fundamental solution would be a high-resolution video generator.
5. Conclusion
In conclusion, we demonstrate that 360° video capture is an effective and practical representation for scene-level 3D generation, as validated by WorldPrompter, a generative pipeline capable of producing navigable 3D scenes from text prompts. By leveraging 360° panoramic video generation as an intermediate representation, combined with a fast 3D reconstruction module, the pipeline enables the synthesis of immersive environments with comprehensive view coverage and high visual fidelity. This scalable approach effectively bridges the gap between text-to-3D generation and full-scene synthesis. We expect this novel idea highlights a promising research direction and has the potential to inspire future work in the field.
Figure 6. Scene generation results. We present our scene generation results by showing (1) the first 360° image from our generated panoramic video as reference, and (2) novel video renderings (Novel View 1-4) from different spatial locations and viewpoints, demonstrating that our approach effectively synthesizes a traversable, 360° viewable scene. More rendered videos can be found in the supplementary documents.
Figure 7. Comparison of scene generation results. We compare our method with LucidDreamer (Chung et al., 2023), DreamScene360 (Zhou et al., 2024) and LayerPano3D (Yang et al., 2024) for text-based scene generation. LucidDreamer generates scenes progressively from a perspective images, synthesizing views one at a time. DreamScene360 and LayerPano3D both use a single panorama as input to an image diffusion model, shown in the first column for reference. For our method, we present the first frame of the generated 360° panoramic video as the reference image. While prior methods can produce visually reasonable results at the initial viewpoint, their range of motion is limited. When synthesizing novel viewpoints (Novel View 1–4) that deviate significantly from the starting position, their reconstructed Gaussian Splatting (GS) scenes exhibit severe artifacts, including blurriness, missing regions, and distorted geometries. As a result, global structural consistency cannot be held. In contrast, our method produces scenes that remain well-aligned with the input prompt and maintain high visual quality even at viewpoints far from the initial position.